OpenStudy (anonymous):

find the indicated term of the given arithmetic sequence a1=115, d=-11, n=12

OpenStudy (anonymous):

12th term is -6 I believe?

OpenStudy (anonymous):

nice, can you please explain how you got that ? I want to learn it : /

OpenStudy (anonymous):

First, An arithmetic sequence is a sequence where each term is found by adding or subtracting the same value from one term to the next. We call this value "common sum" or "common difference"

OpenStudy (anonymous):

So pretty much we start from a1, which is 115, then subtract 11 since it's (-11) until we get to n which is 12

OpenStudy (anonymous):

Another way to find the number is using this equation, (d) x (n) + (a1 plus or minus (d))

OpenStudy (anonymous):

\[-11* 12 + (115-11)\]

OpenStudy (anonymous):

For the last part of the equation you need to do the opposite of the given sign, so instead of subtracting 11, you add 11

OpenStudy (anonymous):

\[-11*12+(115+11) = -6\]

OpenStudy (anonymous):

Do you somewhat get it?

OpenStudy (anonymous):

yeah i somewhat get it a little. thank you

OpenStudy (anonymous):

Take a1 and add or subtract d depending on the sign, until you reach n. Take 115 and subtract 11 (-11) until the 12th term. 115 would be the firt

OpenStudy (anonymous):

1. 115 2. 115-11 =104 3. 104 -11 =93 4. 93-11 = 82 5. 82 -11 =71 6. 71 -11 = 60 7. 60 -11 = 49 8. 49 -11 = 38 9. 38- 11 = 27 10 27 - 11 = 16 11. 16 - 11 =6 ****12th. 5- 11 = -6

OpenStudy (anonymous):

Does that help a bit more?

OpenStudy (anonymous):

Here is a reference site that can help you out. http://www.basic-mathematics.com/arithmetic-sequence.html

OpenStudy (anonymous):

yay ! okay i get it now it took me a minute but i definitely get it now. thank you very much

OpenStudy (anonymous):

yw